OpenStudy (anonymous):

f(x) = 4x + 6, g(x) = 2x^2 Find (fg)(x)

OpenStudy (anonymous):

@Fall12-13

OpenStudy (anonymous):

Okay. F(x) = 4x+6 G(x) = 2x^2 F*g(x) = F(x)*G(x) = (4x+6)*(2x^2) = 8x^3 + 12x^2

OpenStudy (anonymous):

thanks i have more.

OpenStudy (anonymous):

f(x) = 7x + 7, g(x) = 6x^2 Find (fg)(x)

OpenStudy (anonymous):

F(x) = 7x + 7 G(x) = 6x^2 F*G(x) = F(x)*G(x) = (7x+7)*(6x^2) = 42x^3 + 42x^2

OpenStudy (anonymous):

f(x) = 4x + 7, g(x) = 3x^2 Find (f + g)(x).

OpenStudy (anonymous):

F(x) + G(x) = 4x+7 + 3x^2 = 3x^2 + 4x + 7

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

f(x) = 2x + 6, g(x) = 4x^2 Find (f + g)(x)

OpenStudy (anonymous):

2x+6+4x^2